A molecule has r.m.s speed of 550m/s at temperature of 20°c.what is its r.m.s speed at 80°c


1
Expert's answer
2020-07-14T08:53:07-0400

"v_{rms}=\sqrt{\frac{3RT}{M}}"


At "20\u00b0C(293K)" "v_{rms1}=\sqrt{\frac{3R\cdot293}{M}}"


At "80\u00b0(353K)" "v_{rms2}=\sqrt{\frac{3R\cdot353}{M}}"


"\frac{v_{rms1}}{v_{rms2}}=\sqrt{\frac{293}{353}}\to v_{rms2}=v_{rms1}/\sqrt{\frac{293}{353}}="



"=\frac{550}{0.911}=604m/s"